# Interpretable Evaluation of AI-Generated Content with Language-Grounded Sparse Encoders

### v1

@@ Abstract
While the quality of AI-generated contents, such as synthetic images, has become remarkably high, current evaluation metrics provide only coarse-grained assessments, failing to identify specific strengths and weaknesses that researchers and practitioners need for model selection and development, further limiting the scientific understanding and commercial deployment of these generative models. To address this, we introduce Language-Grounded Sparse Encoders (LanSE), a novel architecture that creates interpretable evaluation metrics by identifying interpretable visual patterns and automatically describing them in natural language. Through large-scale human evaluation (more than 11,000 annotations) and large multimodal model (LMM) based analysis, LanSE demonstrates reliable capabilities to detect interpretable visual patterns in synthetic images with more than 93% accuracy in natural images. LanSE further provides a fine-grained evaluation framework that quantifies four key dimensions of generation quality, prompt match, visual realism, physical plausibility, and content diversity. LanSE reveals nuanced model differences invisible to existing metrics, for instance, FLUX’s superior physical plausibility and SDXL-medium’s strong content diversity, while aligning with human judgments. By bridging interpretability with practical evaluation needs, LanSE offers all users of generative AI models a powerful tool for model selection, quality control of synthetic content, and model improvement. These capabilities directly address the need for public confidence and safety in AI-generated content, both critical for the future of generative AI applications.
@@ Introduction
AI-synthesized contents created by generative machine learning models is increasingly used in various aspects of society, including artistic design, advertisement generation, and video production. While recent generative models are remarkably fluent in producing highly realistic outputs [1, 2], they are still prone to notable failure modes, such as hallucinating nonexistent objects [3], misusing lighting cues [4], and violating anatomical coherence [5]. When AI trains on AI-generated content, these failures cascade through systems; for example, medical AI learning from synthetic images with anatomical errors perpetuates those mistakes in real diagnoses [6]. Therefore,it is vital to measure quality of data generated by AI, identify and remove these failures before releasing them for applications and to the public. However, there exists a significant gap in evaluating the quality of generated content and its alignment with human intent. While methods for AI-synthesized data evaluation, typically based on semantic similarity [7], distributional alignment [8, 9], or aesthetics [10], can capture certain aspects of quality, they lack the granularity and interpretability needed to provide nuanced assessments on varied model performance axis [11, 12]. These significant gaps in evaluation can mislead model selection and mask critical failure modes, where models with higher metric scores can still generate outputs with unacceptable structural errors or physics violations [13, 14]. Recent analyses have shown that the fundamental problem of current evaluation approaches is they directly utilize high-dimensional embedding spaces that contain biased information and can ignore critical visual patterns [15, 16].
Unlike many existing evaluation methods that conduct a holistic assessment and assign a single score to each generated output, humans evaluate visual content by identifying distinct visual patterns and considering them collectively [17, 18, 19, 20, 21, 22]. Consider how art critics work: when evaluating a painting created from a prompt, they rarely assign a single score. Instead, they describe which aspects work and which fail: the lighting may be striking, the composition balanced, the anatomy flawed. Previous studies in human cognition have shown that people rely on specific visual cues when judging image quality—such as facial asymmetry [23], inconsistent lighting or shadows [24], and unnatural textures [25]. However, such factors are typically overlooked by existing evaluation metrics, which reduce assessment to single scores or opaque embeddings. This disconnect highlights the urgent need for evaluation methods grounded in interpretable visual patterns, better aligned with how humans assess visual content.
To automatically identify such interpretable visual patterns needed for AI-synthesized image evaluation, recent advances in mechanistic interpretability [26, 27] offer useful tools. Studies have shown that enforcing sparsity on various neural modules elicits monosemantic neurons—individual units that only respond to certain specific concepts [28]. In detail, the linear representation hypothesis [29] reveals that these meaningful features are often encoded as directions in representation space, and modules like sparse autoencoders [30] and transcoders [31] can effectively extract them. Crucially, recent works have found that such interpretable features often correlate with specific failure patterns in generative models [32]. This motivates us to adapt the interpretability modules for evaluation, creating metrics that are both rigorous in computation and interpretable in language to close the gap in evaluating AI generated contents.
Figure 1: Language-Grounded Sparse Encoders (LanSE). Given an input image-caption pair, LanSE maps visual and textual inputs into a sparse, interpretable feature space that reflect evaluation dimensions. LanSE neurons are categorized according to their interpretable visual patterns. Only a few neurons are activated per input.
To solve the problems above, we introduce Language-Grounded Sparse Encoders (LanSE), a method that automatically learn thousands of interpretable visual patterns from training data and use these visual patterns to conduct fine-grained evaluation of synthetic images produced by generative machine learning models (Figure 1). By automatically collecting and concatenating neurons from trained interpretability modules, LanSE filters out noisy neurons, and keeps only the most reliable neurons that serve as high-accuracy visual quality indicators. The semantic meaning of each neuron, which we refer to as an interpretable visual pattern, is automatically identified and phrased by large multimodal models (LMMs). Through extensive human evaluation involving over 11,000 annotations from four independent annotators, LanSE demonstrates a reliable ability to detect nuanced visual patterns, achieving over 93% agreement with human judgments. In addition to natural images, we extended our evaluation pipeline to medical imaging.
Leveraging these interpretable features, LanSE provides a comprehensive evaluation framework that assesses generative models along four critical dimensions: prompt match (whether concepts in prompts exist in images [33]), visual realism (photorealism of generated content [34]), physical plausibility (adherence to intuitive physics constraints [35]), and content diversity (variation in generated outputs [36]). LanSE metrics decompose generative model quality into interpretable dimensions, revealing critical aspects that traditional existing metrics fail to capture or even contradict. To our best knowledge, LanSE provides the first automatic evaluation method for intuitive physics violations, the first semantic alignment metrics and content diversity metrics that support assessments on targeted semantic groups.
By bridging mechanistic interpretability with practical evaluation needs, LanSE improves how we assess generative models today and lays the foundation for more trustworthy, controllable, and safer systems. Its LMM-driven automated pipeline scales the curation of interpretable visual patterns beyond what manual annotation can support. Beyond image generation, LanSE offers a generalizable, interpretable evaluation framework for video synthesis, reinforcement learning, and autonomous-driving perception. The same capacity to extract fine-grained, language-grounded features translates to scientific domains, biomedical imaging, remote sensing, and physics-based simulation, where rigorous quality control is essential. By making model behavior more explainable and reliable, LanSE can strengthen public confidence and safety in AI across education, healthcare, and digital media.
@@ Results
We introduce Language-Grounded Sparse Encoders (LanSE), a general framework for evaluating visual generative models. LanSE detects and names specific visual patterns, such as objects, textures, or structural errors, in synthetic images, whether produced from a text prompt, other inputs, or without any condition. It identifies more than 5,000 such patterns with high reliability, showing strong agreement with over 10,000 human annotations from four independent annotators. From these patterns, we derive four detailed evaluation measures: prompt match (faithfulness to the conditioned description), visual realism (photographic quality), physical plausibility (adherence to the rules of physics), and content diversity (variety of outputs). These measures were validated against more than 3,000 additional human annotations, showing high consistency with human judgment. We then use LanSE to benchmark eight recent generative models of different designs and sizes, revealing clear strengths, weaknesses, and similarities between models—insights that can guide model development and selection. Finally, we compare LanSE with common existing metrics such as FID, CLIP score, and Inception score, showing that LanSE can detect important quality differences missed by these methods.
Figure 2: LanSE neurons and their Human/LMM evaluations. For each of the neuron groups, we randomly sampled 1,000 samples from all the input-neuron pairs with the inputs able to activate the neuron. A total of more than 11,000 human annotations from four independent human annotators, along with 18,000 annotations from two LMMs were collected to determine LanSE’s accuracies.
LanSE Automatically Detects and Names Five Thousand Visual Patterns in Synthetic Images
Modern generative machine learning models learns fine-grained information in a special space within the model encoding interpretable patterns [37, 38]. Various interpretability methods, such as sparse autoencoders [39, 40], transcoders [41], and crosscoders [42] have been introduced to identify these interpretable patterns [43]. However, without careful curation, these identified patterns are often noisy and activating on unrelated patterns and undermining their interpretability [44, 45].
To address this challenge, we develop LanSE, which finds visual patterns in images generated by machine learning models. It does this by collecting millions of candidate neurons from interpretation modules, using LMMs to describe the patterns each neuron detects, and keeping only those with high accuracy (Figure 2).
LanSE uses LMMs to systematically curate neurons at scale. It begins by aggregating neurons from 250 interpretability modules trained on diverse datasets, producing 3.75 million candidates. In a two-stage process, LMMs first examine each neuron’s most strongly activated images to generate a natural language description of the visual pattern it detects. Neurons are then filtered by their measured accuracy, keeping only those that match their described pattern more than 80% of the time (see Method).
@@
This process yields 5,309 high-quality neurons, each corresponding to a distinct visual pattern found in synthetic images. These patterns are grouped into nine categories: semantic content—(1) human, (2) animal, (3) object, (4) activity, (5) environment; unrealistic patterns—(6) style, (7) artifact; and physics violations—(8) distortion, (9) structure. Each category plays a specific role in LanSE’s evaluation framework (examples in Appendix A).
Visual Patterns Identified by LanSE Align with Human Annotations
@@
To understand if the visual patterns detected by LanSE are both accurate and reliable for evaluating image quality, we validated them against large-scale human and automated assessments. We conducted evaluations involving over 11,160 human annotations from four independent annotators (see Appendix F). Each human annotator reviewed 2,790 neuron–image pairs, where each pair comprised a neuron’s interpreted description and an image that activated it, and judged whether the activation matched the described pattern. As shown in Figure 2, LanSE neurons achieved 93% accuracy under human evaluation, with category-specific performance ranging from 76% to 99%. The four annotators achieved an inter-annotator agreement of 85.93%. In addition to human annotation, we also included automated assessments from two state-of-the-art large multimodal models (LMMs). On the same dataset, automated evaluation of LanSE show strong agreement with human annotation achieving 90.4% accuracy with category-specific performance ranging from 77.3% to 98.7%.
Derivation of Diagnostic Metrics for Evaluating Synthetic Images Using Visual Patterns Identified by LanSE
@@
Having introduced LanSE and validated the quality of its detected visual patterns, we now show how these patterns induce four diagnostic metrics for synthetic-image quality: prompt match, visual realism, physical plausibility, and content diversity (See Appendix E). These metrics target complementary aspects that common aggregate measures fail to distinguish. Figure 3 summarizes the computations.
@@
Formally, let $\mathrm{LanSE}(x,y)=\boldsymbol{n}=[n_{1},\ldots,n_{d}]\in\{0,1\}^{d}$ denote the sparse activation vector for an image–text pair $(x,y)$. Each neuron $n_{i}$ is associated with a natural-language explanation $\epsilon_{i}$ (e.g., “horses and equestrian-related activities”). The $d$ neurons are partitioned into 3 disjoint groups ($\mathcal{C}={\mathbf{prompt},\mathbf{real},\mathbf{phy}}$). For $g\in\mathcal{C}$, we write $\mathrm{N}_{g}(x,y)\in\{0,1\}^{d_{g}}$ for the subvector of activations in group $g$.
Figure 3: Evaluation metrics based on LanSE neurons. We define four evaluation metrics utilizing the activations of LanSE’s interpretable neuron groups. They each measure a distinct aspect of generative model’s performances.
@@
Diagnostic Metrics from LanSE Align with Human Judgments
To assess whether the diagnostic metrics automatically derived from LanSE meaningfully capture quality distinctions in synthetic images, we performed meta-evaluations comparing each metric against human judgments. For the first three metrics—prompt match, visual realism, and physical plausibility—we constructed dedicated validation sets and labeled them using human annotators. Human evaluation of content diversity requires nuanced expertise to assess differences across output distributions and was therefore excluded from this validation.
The validation protocol was as follows: from a pool of 3,000 generated images, two independent annotators classified each image for each metric category as positive (containing the corresponding error type) or negative (error-free). An image was assigned to the positive set if the annotator indicated the presence of the relevant error. We then computed the average metric scores for the positive and negative sets to assess discriminative power. Our results shown in Figure 4 demonstrate strong alignment between LanSE metrics and human judgments. Across all three metrics, images in the negative sets show significantly lower diagnostic metrics than those in the positive sets.
@@
Evaluation of Generative Models Using Diagnostic Metrics from LanSE
@@
Our results presented in Figure  5 reveal distinct performance profiles for each model. FLUX achieves the highest physical plausibility scores, making it particularly suitable for applications requiring strict physical plausibility. SDXL-medium excels in both photorealism and content diversity, producing the most realistic and varied outputs. DALL·E 3 demonstrates superior prompt match but at the cost of reduced output diversity. Interestingly, we observe that generated images generally exhibit stronger semantic alignment with their captions than real-world image-caption pairs, suggesting that current models may be over-optimized for prompt match. Furthermore, our results reveal a clear correlation between model scale and physical plausibility, larger models significantly reduce distortion and structure errors. Figure 6 presents detailed group-wise performance metrics that enable more nuanced model selection. These fine-grained insights provide practitioners with actionable guidance for choosing models based on their specific application requirements, whether prioritizing prompt match, visual realism, physical plausibility or content diversity.
@@
Figure 5: Evaluation results for generative models. Performance comparison of eight generative models across four LanSE metrics: prompt match, visual realism, physical plausibility, and content diversity. Lower scores indicate better performance for the first three metrics, while higher scores indicate better content diversity.
@@
Figure 6: Model leaderboards for group-wise scores. Different generative models exhibit distinct performance profiles across evaluation dimensions. For prompt match, environmental elements represent the most common source of semantic misalignment across all models. FLUX achieves superior performance in human and object semantics, while SDXL-medium excels at animal semantics, DALL·E3 at activity semantics, and SDXL-turbo at environmental semantics. In visual realism metrics, SDXL-medium demonstrates the strongest performance in avoiding both unmentioned art styles and invented visual artifacts, with SDXL-large as the second-best performer. Physical plausibility metrics reveal FLUX as the leading model in preventing both surface distortions and anatomically implausible structures, followed by SDXL-medium and SDXL-large. Content diversity analysis shows complementary strengths: SDXL-large produces the most varied human representations, SDXL-medium generates the greatest diversity in animals, objects, and activities, while DALL·E3 achieves the highest environmental variety, and SDXL-large gives the best style diversity.
Inter-model Correlation Analysis via LanSE
@@
To examine similarity and difference between generative models at the level of neural activation patterns, we analyze the consistency with which different models trigger LanSE neurons. For each model, we generate 6,000 image–caption pairs and compute pairwise correlation matrices of LanSE activations across all nine neuron groups.
@@
Figure 7 reveals two main findings. First, content neurons, covering human, animal, object, activity, environment, and style, exhibit high inter-model correlations (mean $r=0.74$), whereas error-signaling neurons, artifact, distortion, and structure, show markedly lower correlations (mean $r=0.31$). This suggests that models tend to converge on semantic representations but diverge substantially in their failure modes. Second, FLUX and SDXL-Medium display notably higher correlations, particularly in style ($r=0.60$), environment ($r=0.82$), and activity ($r=0.76$) neurons, implying shared architectural or training strategies. By contrast, SDXL-Turbo shows the most distinctive activation patterns, with the lowest average style correlation ($r=0.51$).
Figure 7: Conducting correlation analysis on generative models with LanSE. Pairwise correlation matrices computed from LanSE neuron activations across six generative models, with each matrix representing one of nine neuron groups. Warmer colors indicate higher correlation between model pairs.
Comparison of LanSE Metrics with Existing Metrics
To position LanSE within the landscape of generative model evaluation, we compare our metrics with widely-adopted baseline methods: Fréchet Inception Distance (FID) [8], Inception Score (IS) [9], and CLIP Score [59]. Figure 8 reveals several key insights. First, existing metrics show limited discriminative power among recent models. CLIP scores cluster tightly (0.269-0.285), failing to capture the substantial quality differences we observe through LanSE. FID scores, while showing more variation, contradict human preferences, SDXL-base achieves the best FID (6.69) despite producing lower-quality outputs than FLUX or SDXL-medium. This suggests FID’s distributional matching may not align with perceptual quality.
Figure 8: LanSE reveals model differences that traditional metrics fail to capture. Comparatively, evaluation results across traditional metrics show limited discriminative power and occasional misalignment with human preferences.
To understand these limitations, we conduct detailed analysis comparing CLIP scores with our prompt match metric. While CLIP is widely used for semantic alignment evaluation, it exhibits systematic biases. Figure 9 reveals a critical pattern: image-text pairs with high prompt match errors never achieve high CLIP scores (empty upper-right regions), indicating that severe misalignments are detectable by both metrics. However, the scatter plots show substantial spread in the lower-left regions, where CLIP assigns low scores to images that LanSE identifies as semantically aligned in specific aspects.
This discrepancy stems from CLIP’s bias toward object existence while overlooking compositional details [16]. Figure 10 demonstrates this limitation through examples where CLIP assigns high scores (>0.28>0.28) to images with clear semantic mismatches, unmentioned objects, incorrect activities, or wrong environmental contexts, that LanSE successfully identifies through its specialized neuron groups.
These findings demonstrate that while existing metrics provide useful aggregate measures, they lack the granularity and interpretability needed for actionable model improvement. LanSE’s decomposed evaluation along specific visual patterns enables practitioners to understand not just that models differ, but precisely how and why they differ, critical information for both model selection and development.
Figure 9: Clip v.s. $\mathcal{M}_{prompt}$. Each point represents an image–caption pair. Notably, regions with both high CLIP-T and high $\mathcal{M}_{prompt}$ are often sparse or unpopulated.

### v2

@@ Abstract
The rapid development of generative AI has transformed content creation, communication, and human development. However, this technology raises profound concerns in high-stakes domains, demanding rigorous methods to analyze and evaluate AI-generated content. While existing analytic methods often treat images as indivisible wholes, real-world AI failures generally manifest as specific visual patterns that can evade holistic detection and suit more granular and decomposed analysis. Here we introduce a content analysis tool, Language-Grounded Sparse Encoders (LanSE), which decompose images into interpretable visual patterns with natural language descriptions. Utilizing interpretability modules and large multimodal models, LanSE can automatically identify visual patterns within data modalities. Our method discovers more than 5,000 visual patterns with 93% human agreement, provides decomposed evaluation outperforming existing methods, establishes the first systematic evaluation of physical plausibility, and extends to medical imaging settings. Our method’s capability to extract language-grounded patterns can be naturally adapted to numerous fields, including biology and geography, as well as other data modalities such as protein structures and time series, thereby advancing content analysis for generative AI.
@@ Introduction
For millennia, human-created content served as the exclusive record of civilization, with generally every painting, text, and diagram traceable to human thought and effort [1, 2, 3]. Generative AI has brought an end to this era [4], producing seemingly indistinguishable alternatives at unprecedented scale [5, 6, 7], permeating every domain from scientific publications [8, 9] to medical diagnostics [10, 11] to artistic designs [12, 13]. Yet this transformative technology raises profound concerns about its deployment across critical domains: from diagnostic reliability in medical applications [14, 15, 16] and accelerated misinformation propagation [17, 18] to cultural homogenization [19, 20, 21] and data contamination [22, 23]. These consequential risks necessitate immediate development of rigorous methods to analyze, evaluate, and control AI-generated content, protecting the integrity and trustworthiness of AI systems in high-stakes applications.
Content analysis for generative AI encounters a fundamental scope-specificity dilemma. While effective content analysis requires comprehensive assessment across multiple quality dimensions, from prompt adherence to anatomical correctness, real-world failures typically manifest as specific visual patterns that require precisely the granular detection capabilities that many broad-scope methods fail to provide [24]. Existing analytic methods typically prioritize either comprehensive scope or detection specificity. Some analytic approaches treat generated content as indivisible units, providing holistic assessments through distributional alignment metrics [25, 26] or semantic similarity scores [27, 28], while other developments invent specialized detection tools for specific failure modes—such as text generation errors [29, 30], model hallucination [31], or anatomical anomalies [32]. These specialized approaches achieve high precision but lack adaptability across different content domains.
Comparatively, human evaluation of AI-generated content operates through fundamentally different mechanisms than automated analytic approaches. Consider how art critics assess AI-generated paintings: rather than providing holistic scores, they articulate which aspects work and which fail: the lighting may be striking, the composition balanced, the anatomy flawed, which are all inherently visual patterns interpreted with natural language. This pattern-based evaluation process reflects broader cognitive mechanisms documented in human visual assessment [33, 34, 35, 36, 37, 38, 39, 40], suggesting a promising approach to content analysis: developing methods that identify language-grounded visual patterns in AI-generated content.
Here we introduce Language-Grounded Sparse Encoders (LanSE), a content analysis method that systematically identifies interpretable visual patterns in AI-generated content and grounds them with natural language descriptions. LanSE leverages recent advances in mechanistic interpretability [41, 42], particularly sparse autoencoders [43, 44, 45] and transcoders [46, 47, 48], combined with large multimodal models (LMM) analysis to comprehensively discover, interpret and organize visual patterns across multiple content aspects. We validate LanSE’s content analysis capability through extensive human evaluation involving over 11,000 annotations from six independent annotators. This evaluation demonstrates strong alignment with human judgment, achieving 93% agreement when detecting visual patterns in natural images. This robust validation establishes LanSE as a reliable foundation for systematic content analysis across diverse applications.
LanSE’s core capability to comprehensively identify visual patterns enables three complementary applications in content analysis for generative AI. For generative model evaluation, LanSE enables systematic assessment through four evaluation metrics: prompt match (whether concepts in prompts exist in the output [49]), visual realism (whether the output is photorealistic [50]), physical plausibility (whether the output adheres to physics constraints [51]), and content diversity (whether outputs show sufficient variation [52]). LanSE enables systematic automated evaluation of physical plausibility in AI-generated content and reveals failure modes that existing methods overlook. For content understanding, LanSE provides interpretable multi-label annotations that enable natural language-based image retrieval, systematic dataset curation, and granular failure mode detection across thousands of visual patterns. Beyond single-model analysis, LanSE enables inter-model correlation analysis that reveals how different generative systems encode visual knowledge, showing that models converge on semantic representations while diverging in their specific failure modes. For cross-domain analysis, the core methodology extends naturally beyond natural images: we demonstrate clinical pattern detection in medical images achieving 74% agreement with expert radiologists, with potential applications spanning satellite imagery analysis, biological microscopy analysis, and scientific data interpretation. We believe LanSE represents a critical step toward transparent and accountable AI systems, addresses the content analysis need for this generative AI era, and paves the way to interpretable AI architectures where human understanding guides system behavior.
Figure 1: Content Analysis with Language-Grounded Sparse Encoders. LanSE systematically identifies visual patterns in AI-generated content and describes them in natural language (e.g., ”dogs in outdoor environments” in natural images and ”evidence of bibasilar atelectasis” in medical images). These patterns enable systematic content analysis across multiple applications: quality evaluation, multi-label annotation, and inter-model analysis of generative models across various domains.
@@ Results
LanSE identifies 5,309 reliable visual patterns across natural images, validated through 11,160 human assessments achieving 93% agreement on pattern-image correspondence. We extend LanSE to medical imaging, building CXR-LanSE that identifies 899 diagnosis-relevant patterns with 74% alignment with expert radiologists. For individual content analysis, LanSE decomposes images into interpretable visual patterns, enabling fine-grained understanding of content properties and systematic dataset characterization. For generative model analysis, we design four diagnostic metrics that enable decomposed analysis of critical aspects: prompt match, visual realism, physical plausibility, and content diversity. Using these metrics, we evaluate eight state-of-the-art generative models and reveal performance characteristics that existing metrics fail to capture. Furthermore, we conduct inter-model correlation analysis revealing that generative models converge in semantics but diverge in failure modes.
LanSE Identifies 5,309 Visual Patterns with 93% Accuracy
@@
LanSE systematically discovers visual patterns by collecting millions of candidate neurons from interpretability modules, using LMMs to describe the visual patterns each neuron represents, and retaining only those with accuracies greater than 80% (detailed in Method). This process yields 5,309 high-quality neurons, each corresponding to a distinct visual pattern found in AI-generated images. These patterns are grouped into three main categories with nine subgroups: semantic content—(1) human, (2) animal, (3) object, (4) activity, (5) environment; unrealistic patterns—(6) style, (7) artifact; and physics violations—(8) distortion, (9) structure. Each category plays a specific role in LanSE’s evaluation method (examples in Appendix A).
@@
To understand if the visual patterns detected by LanSE are both accurate and reliable for evaluating image quality, we validated them against large-scale human and automated assessments. We conducted evaluations involving over 11,160 human annotations from four independent annotators (see Appendix F). Each human annotator reviewed 2,790 neuron–image pairs, where each pair comprised a neuron’s interpreted description and an image that activated it, and judged whether the activation matched the described pattern. As shown in Figure 2, visual patterns discovered by LanSE achieved 93.2% accuracy under human evaluation, with category-specific performance ranging from 86.0% to 99.9%. The four annotators achieved an inter-annotator agreement of 85.93%. In addition to human annotation, we also included automated assessments from two state-of-the-art LMMs. On the same dataset, automated evaluation via LMM shows strong agreement with human annotation achieving 90.4% accuracy with category-specific performance ranging from 76.2% to 98.5%.
CXR-LanSE Detects 899 Clinical Relevant Visual Patterns
CXR-LanSE extends our content analysis approach to medical imaging by systematically discovering clinical visual patterns in chest X-ray images. Using the same interpretability framework adapted for medical data, CXR-LanSE identifies 899 distinct radiological patterns from MIMIC-CXR datasets and AI-generated medical images. These patterns capture clinically meaningful features ranging from anatomical landmarks to pathological findings and medical devices. To validate the clinical relevance of discovered patterns, we conducted evaluations with two expert radiologists who assessed pattern-image correspondence on randomly selected neuron-image pairs. Each pair consisted of a CXR-LanSE pattern description and a chest X-ray image that activated the corresponding neuron. The radiologists judged whether the automated interpretation accurately captured clinically meaningful features present in the image.
CXR-LanSE achieved 74% agreement with expert radiologists, with representative patterns including ”evidence for bibasilar atelectasis,” ”evidence for pulmonary opacities,” and ”evidence for central venous catheters.” This validation demonstrates CXR-LanSE’s capability to automatically identify clinically relevant visual patterns without requiring manual feature engineering or domain-specific training beyond the underlying multimodal representations. The systematic discovery of interpretable medical patterns enables applications in clinical decision support, automated radiology quality assessment, and AI-generated medical image validation, addressing critical needs for explainable AI in healthcare settings.
Figure 2: Visual patterns identified by LanSE for natural and medical images. A total of 5,309 different visual patterns in natural images and 899 visual patterns in medical images are automatically identified by LanSE. More than 11,000 human annotations from in total six independent human annotators, along with 18,000 annotations from two LMMs are collected to determine the accuracy of these visual patterns.
LanSE Enables Pattern-based Datapoint Analysis
LanSE provides detailed multi-label annotations for individual images by identifying all activated visual patterns, enabling comprehensive content characterization at the datapoint level as demonstrated in Figure 3. Across our test set of 48,000 generated images, we observe that images typically activate 15-30 of the 5,309 LanSE neurons. Individual images reveal complex pattern combinations—for instance, simultaneously activating neurons for ”depiction of human figures” (human category), ”transportation-related elements in urban settings” (environment category), and ”anatomically impossible structures” (structure category), providing comprehensive characterization of image samples.
This multi-labeling capability enables transformative applications across domains. First, LanSE enables natural language-based image retrieval, allowing users to search datasets using semantic queries that combine multiple visual patterns. Second, integrating LanSE into training loops could enable pattern-aware optimization, where models receive granular feedback about specific visual patterns they fail to generate correctly. Third, CXR-LanSE enables clinical interpretability by automatically detecting and describing radiological features in natural language.
Figure 3: Multi-label pattern detection enables detailed image analysis across domains. LanSE automatically identifies all activated visual patterns in generated images (top) and medical images (bottom), providing comprehensive semantic fingerprints that characterize both content and potential errors. This granular annotation supports applications from dataset curation to targeted model debugging.
LanSE Provides Diagnostic Metrics for Model Analysis
@@
LanSE’s comprehensive visual pattern vocabulary enables the construction of four diagnostic metrics that systematically characterize different aspects of generated content: prompt match, visual realism, physical plausibility, and content diversity (detailed definition in Appendix E). These metrics target complementary aspects that various existing metrics fail to distinguish as shown in Figure 4.
@@
Formally, let $\mathrm{LanSE}(x,y)=\boldsymbol{n}=[n_{1},\ldots,n_{d}]\in\{0,1\}^{d}$ denote the sparse activation vector for an image–text pair $(x,y)$. Each neuron $n_{i}$ is associated with a natural-language explanation $\epsilon_{i}$ (e.g., “horses and equestrian-related activities”). The $d$ neurons are partitioned into 3 disjoint groups ($\mathcal{C}=\{\mathbf{prompt},\mathbf{real},\mathbf{phy}\}$). For $g\in\mathcal{C}$, we write $\mathrm{N}_{g}(x,y)\in\{0,1\}^{d_{g}}$ for the subvector of activations in group $g$.
Figure 4: Image diagnostic metrics derived from LanSE visual patterns. Utilizing visual patterns obtained using LanSE, we define four diagnostic metrics for AI-generated natural images, each measuring a distinct aspect of the generative model.
@@
Diagnostic Metrics Align with Human Quality Judgments
To assess whether the diagnostic metrics automatically derived from LanSE validly capture quality distinctions in AI-generated images, we performed meta-evaluations comparing each metric against human judgments. For the first three metrics—prompt match, visual realism, and physical plausibility—we construct dedicated validation sets and label them using human annotators. Human evaluation of content diversity requires nuanced expertise to assess differences across output distributions and was therefore excluded from this validation.
The validation protocol was as follows: from a pool of 3,000 generated images, two independent annotators classified each image for each metric category as positive (containing the corresponding error type) or negative (error-free). An image was assigned to the positive set if the annotator indicated the presence of relevant error. We then compute the average metric scores for the positive and negative sets to assess discriminative power. Our results shown in Figure 5 demonstrate a strong alignment between LanSE metrics and human judgments. Across the three metrics, the images in the negative sets show significantly lower metric values than those in the positive sets.
@@
Fine-grained Evaluation of Eight Prevalent Generative Models
@@
Our results presented in Figure 6 reveal distinct performance profiles for each model. FLUX achieves the highest physical plausibility scores, making it particularly suitable for applications requiring strict physical plausibility. SDXL-medium excels in both photorealism and content diversity, producing the most realistic and varied outputs. Figure 7 presents detailed group-wise performance metrics that reveal more nuanced characteristics. These fine-grained insights provide practitioners with actionable guidance for choosing models based on their application requirements, whether prioritizing prompt match, visual realism, physical plausibility or content diversity.
In the evaluation of recent generative models, our results reveal a striking pattern: while they achieve strong prompt match, often exceeding that of real-world datasets, they consistently struggle with visual realism, physical plausibility, and content diversity. Recent models frequently exhibit distinctive stylistic signatures that deviate from photorealistic representation, generate distorted surfaces, and produce physically implausible structures. Their content diversity remains significantly below that of natural image distributions. These findings suggest the machine learning community should recalibrate its priorities, shifting focus from prompt match, a largely solved problem, toward these more nuanced aspects of visual generation.
@@
Figure 6: Evaluating generative machine learning models with LanSE-derived diagnostic metrics. Performance comparison of eight generative machine learning models across four LanSE-derived diagnostic metrics: prompt match, visual realism, physical plausibility, and content diversity. Lower scores indicate better performance for the first three metrics, while higher scores indicate better content diversity.
@@
Figure 7: Model leaderboards for subgroup scores of LanSE-derived diagnostic metrics. Different generative models exhibit distinct performance profiles across evaluation dimensions. For prompt match, environmental elements represent the most common source of semantic misalignment across all models. FLUX achieves superior performance in human and object semantics, while SDXL-medium excels at animal semantics, DALL·E3 at activity semantics, and SDXL-turbo at environmental semantics. In visual realism metrics, SDXL-medium demonstrates the strongest performance in avoiding both unmentioned art styles and invented visual artifacts, with SDXL-large as the second-best performer. Physical plausibility metrics reveal FLUX as the leading model in preventing both surface distortions and anatomically implausible structures, followed by SDXL-medium and SDXL-large. Content diversity analysis shows complementary strengths: SDXL-large produces the most varied human representations, SDXL-medium generates the greatest diversity in animals, objects, and activities, while DALL·E3 the highest in environment and SDXL-large the highest in style.
LanSE Enables Inter-model Correlation Analysis
@@
LanSE’s visual patterns enable correlation analysis that reveals fundamental similarities and differences between generative models. We generate 6,000 image-caption pairs for each model and compute pairwise correlation matrices across all neuron groups.
@@
Figure 8 reveals two striking patterns in model behavior. First, content neurons, covering human, animal, object, activity, environment, and style, exhibit high inter-model correlations (mean $r=0.74$), whereas error-signaling neurons, artifact, distortion, and structure, show markedly lower correlations (mean $r=0.31$). This demonstrates that while models converge on semantic representations, they develop distinct failure modes. Second, specific model pairs show notable similarities: FLUX and SDXL-medium display particularly high correlations in style ($r=0.60$), environment ($r=0.82$), and activity ($r=0.76$) neurons, suggesting shared architectural or training strategies. Conversely, SDXL-turbo exhibits the most distinctive activation patterns, with the lowest average style correlation ($r=0.51$).
Figure 8: Generative models share semantic understanding but produce unique errors. Pairwise correlation matrices of LanSE visual patterns across six models reveal high inter-model agreement for content-related neurons but substantial divergence in error-detecting neurons. Warmer colors indicate higher correlations.
LanSE Reveals Model Differences Missed by Prevalent Methods
To demonstrate LanSE’s advantages for content analysis, we compare our diagnostic capabilities with widely-adopted baseline methods: Fréchet Inception Distance (FID) [25], Inception Score (IS) [26], and CLIP score [27].
Our analysis aligns with CLIP’s systematic biases identified in various previous works [67]. While both CLIP and our prompt match metric generally detect severe misalignments (empty upper-right regions in Figure 10), CLIP can assign high scores to outputs with obvious semantic mismatches when partial object alignment masks compositional errors, as shown in Figure 9. In contrast, LanSE’s prompt match evaluation is based on visual patterns and can capture subtle semantic mismatches CLIP fails to detect.
Figure 11 further demonstrates two critical limitations of these prevalently used metrics. First, they show minimal discriminative power—CLIP scores cluster tightly (0.269-0.285) despite substantial quality differences between models. Second, they misalign with human perception; SDXL-base achieves the best FID and Inception Score while producing notably lower-quality outputs than FLUX or SDXL-medium, suggesting these metric values poorly reflect perceptual quality.